OpenStudy (anonymous):

How do you factor this equation x^2-5^2

OpenStudy (anonymous):

\(a^2-b^2=(a+b)(a-b)\) here, a=x b=5

OpenStudy (anonymous):

how did you get to this answer?

OpenStudy (anonymous):

thats the general formula that u can use.

OpenStudy (anonymous):

no, it would be (x-5)(x+5)

OpenStudy (anonymous):

ohhh i see so thats why ive been getting the answers wrong lol

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

welcome ^_^

OpenStudy (anonymous):

wait just to make sure i the equation would end up like this right? (x+5)(x-5)

OpenStudy (anonymous):

exactly.

OpenStudy (anonymous):

but how come the answer ends with the signs switched?

OpenStudy (anonymous):

because the formulas say so if you multiply them out, (x+5)(x-5) = x*x -x*5 +5*x -5*5 5x gets cancelled, and you are left with x^2-5^2

OpenStudy (anonymous):

ahh so for every problem would i have to switch the signs?

OpenStudy (anonymous):

only when its difference of squares, like a^2-b^2

OpenStudy (anonymous):

ahh alright so if the equation was like 2x-25 you wouldnt switch it you would leave the answer as (x+5)(x-5)

OpenStudy (anonymous):

2x-25 is not of the type a^2-b^2

OpenStudy (anonymous):

ahh my bad 5^2-6^2 like the squares have to be the same in order for it to be left alone

OpenStudy (anonymous):

5^2-6^2 will then be (5+6)(5-6)

OpenStudy (anonymous):

thank you very much!! you helped me out big time :P

OpenStudy (anonymous):

most welcome ^_^